Given : n = 6, ∑(x − 18.5) = −3, ∑(y − 50) = 20, ∑(x − 18.5)(y − 50) = −120,

∑(x − 18.5)

2 = 19, ∑(y − 50)

2 = 850 . Calculate coefficient of correlation.


1
Expert's answer
2021-05-24T16:25:59-0400

Given


"n=6, \bar{x}=18.5, \bar{y}=50"

"S_{xx}=\sum(x-\bar{x})^2=19"

"S_{yy}=\sum(y-\bar{y})^2=850"

"S_{xy}=\sum(x-\bar{x})(y-\bar{y})=-120"

Coefficient of correlation


"r=\dfrac{S_{xy}}{\sqrt{S_{xx}}\sqrt{S_{yy}}}"

"=\dfrac{-120}{\sqrt{19}\sqrt{850}}\approx-0.944267"

"r<0, 0.7<|r|\leq1"

Strong negative correlation.